# Scaling Laws for Optimal Data Mixtures

### v1

@@ 7 Related works
Scaling laws. Scaling laws research investigates how model performance varies with training compute. Foundational studies [29, 32, 30] established that language models follow a predictable power-law relationship between performance and compute, allowing for the optimal allocation of parameters and training tokens within a specified budget. Scaling behavior has since been explored across a wide range of domains, including vision models [20, 52], diffusion transformers [36], and other fields [12, 48]. While typical scaling laws consider the number of total parameters, other studies have examined the influence of both width and depth [44], or the number of parameters allocated to the teacher and student in cse of model distillation [9]. Sparse Mixture of Experts (MoE) models have been another focus, with investigations into how factors like sparsity, the number of experts, and routing strategies affect scaling [33, 14, 63, 2]. For multimodal models, scaling laws have been explored in studies such as [3, 55]. Of particular relevance is [55], which examines native multimodal models. However, their analysis is constrained by a fixed pretraining mixture.
@@
Scaling laws for data mixtures Optimizing data mixtures for model training is a critical challenge, often requiring extensive experimentation. Recent studies have begun exploring systematic approaches to identify optimal mixtures more efficiently. For instance, Goyal et al. [24] investigated scaling laws for data filtering in CLIP models, emphasizing data quality and repetition. Gu et al. [27] examined scaling laws for continual pretraining of language models, predicting the optimal balance between pretraining and domain-specific data, while Bethune et al. [7] followed the same approach, focused on forgetting in finetuning. Similarly, [11] derived scaling laws that account for data quality factors such as diversity. Closer to our work, Ye et al. [67] and Ge et al. [22] propose scaling laws that model the loss as a function of $h$ for fixed (N, D), but they do not consider a joint law for $(N,D,h)$ as we do here. We also find that, in our experiments, for a fixed (N, D), our scaling laws extrapolate better to unseen mixtures (see appendix B). Further, these approaches are generally constrained to a single modality, and they consider relatively small models, below 1B parameters.
@@
Data mixture selection. The standard approach to selecting training data mixtures relies on trial and error, where different combinations are tested to determine the best-performing mixture [57, 37, 34, 43, 69, 20, 54]. However, this method is costly, leading to recent efforts exploring alternative strategies. Some studies adopt heuristic methods, adjusting mixture ratios based on data sizes for each domain [51, 26, 13] or to match a target task’s distribution [25]. Others predict model performance using small models that take the mixture as input [66, 4, 39, 18]. A third approach employs auxiliary models to rank and select high-quality training data, which has been popularized recently by large foundation models [64, 8, 65, 49, 16].
@@ 8 Discussion
Limitations. Our current study is focused on pretraining, but continual pre-training and finetuning are also scenarios in which the mixture is important. Our scaling law predicts a generic target loss [32], which is known to correlate with downstream task performance [30, 42]. Future work may involve predicting this performance directly, like [31]. Furthermore, assuming no data repetition (i.e., an infinite stream of data from each domain), as is typical for LLM pretraining, is unrealistic when training with very scarce, high-quality sources. Finally, we assume that the mixture is fixed throughout training, but future works may consider a dynamic evolution of the weights (e.g., curriculum learning).
@@
Conclusion. We propose a data mixing law that predicts the loss on an arbitrary target domain, from both mixture coefficients and the compute budget $(N,D)$. Our laws hold for language, multimodal, and vision models pretraining. The optimal mixture coefficients found from a small scale can be used for much larger models and training budgets, demonstrating significant improvement over domain weights found by naive grid search. This work paves the way to a principled theory of data mixture selection.
9 Aknowledgment
We thank Victor Guilherme Turrisi da Costa for technical support. We thank Marco Cuturi, Jason Ramapuram, Tatiana Likhomanenko, Denise Hui, Samy Bengio, and the MLR team at Apple for their support throughout the project.
@@ Appendix A Implementation details
Optimizer Fully decoupled AdamW [40]
@@ A.1 LLM pretraining
For the main experiments, we use LLAMA style architectures. For the analyses with the Pile domains, we borrow architectures from [8]. We use a fixed depth of 24 and change the latent dimension of the network to obtain different model scales. All hyperparameters are described in Tab. 5 and Tab. 7.
@@ A.2 Multimodal pretraining
We closely follow the implementation of [55] and present in Tab. 6 the pre-training hyperparameters for the model configurations used in our scaling laws study. Our models range from 100M to 8B parameters, with width scaling accordingly while maintaining a constant depth of 24 layers. We use causal attention for text tokens and bidirectional attention for image tokens. Learning rates are adjusted based on model size, generally decreasing for larger models. These values were determined through empirical testing. Optimization is handled using a fully decoupled AdamW optimizer with momentum values set to $\beta\_1=0.9$, $\beta\_2=0.95$, and a weight decay of $1\times 10^{-4}$. Each training batch consists of 2,000 samples, totaling 2 million tokens with a 1,024-token context length. Gradients are clipped at 1.0, and training begins with a warmup phase of 1,000 iterations, followed by a constant learning rate schedule to reduce the number of experiments.
For vision inputs, we process images as $(14,14)$ patches with augmentations including Random Resized Crop (224px, scale range of [0.4, 1.0]) and Random Horizontal Flip with a 50% chance. Model training benefits from efficiency techniques such as bfloat16 precision, Fully Sharded Data Parallel (FSDP) [70], activation checkpointing, gradient accumulation, and sequence packing to minimize padding in the image-captioning dataset.
We assess model performance on three held-out data subsets: interleaved data (Obelics), image-caption data (HQITP), and text-only data (DCLM), following prior works [30, 3, 2]. This setup provides a robust evaluation of model generalization across diverse data types.
@@
Optimizer Fully decoupled AdamW [40]
@@
Optimizer AdamW [40]
@@ Appendix B Detailed extrapolation results
Table 9: Full results of experiments in Fig. 3 for the multimodal experiment.
@@
Table 10: Full results of experiments in Fig. 3 for the LVM experiment.
@@
Table 11: Full results of experiments in Sec. 6 for the LLM experiment.
@@
Table 12: Full results of experiments in Fig. 3 for the cosine schedule experiment.
@@
B.1 Comparison to the laws of Ye et al. [67]
Ye et al. [67] propose four laws to model the behavior of the loss on a domain as a function of $h$ only, that is, for a fixed $N,D$ budget. They propose the following formulas, rewritten with notations consistent with our notation:
@@
where the parameters of the law are the $E,C,C_{i},\gamma_{i}$. We compare this with the form of our scaling law eq. 4 when $N$ and $D$ are fixed:
@@
We fit all those scaling laws on the LLM training data, keeping only one value for $N,D$ (we take $N=200m$, $D=8B$ tokens). We only keep $25$ training mixtures to fit the laws, and report the MRE on the $84-25=59$ remaining in Tab. 13. Note that we had trouble fitting the M3 law, which is also reported to underperform in [67]. Overall, our formula gives systematically better training errors, and it most of the time translates to better testing errors on unseen mixtures. We stress once again that one of the core contributions of our work is to explain how scale interacts with data mixtures, by proposing scaling laws that take $N,D$, and $h$ as inputs. [67] only consider scaling laws with respect to $h$.
Table 13: Comparison of our scaling laws for a fixed (N, D) budget with those of [67] on the LLM data.
@@
B.2 Comparison to Ge et al. [22]
Ge et al. [22] propose a scaling law to evaluate the loss the $i-th$ training domains, as a function of both $h$ and number of tokens $D$:
@@
We consider a fixed size of model (N=200M) on the LLM training experiment, take only $25$ mixtures to fit the scaling laws, and report the MRE on the remaining testing set in Tab. 14. We see that our proposed scaling law achieves a significantly lower MRE on both train and testing data, highlighting the importance of taking all other domains into account.
Table 14: Comparison of our scaling laws for a fixed model size $N$ with that of [22] on the LLM data.
@@
Appendix C Additional results
We evaluate 7B LLMs on different downstream tasks, and report the results in Table 15.
Table 15: Evaluations of the 7B models. We report the median score on the CORE tasks [35] as well as other accuracies on standard benchmarks. The model trained with weights that minimize the loss on OpenHermes performs best.
$h^{*}_{avg}$ 56 32 71 40 60 79 52 72 65
$h^{*}_{OH}$ 58 37 70 40 67 79 53 72 65
Uniform 53 30 70 40 46 78 51 70 65
Base 52 25 70 43 51 79 53 71 65
Appendix D Additional analysis
D.1 Optimal domain weights
Table 16: Optimal domain weights $h^{*}$ for LLMs. Note that for the OpenHermes optimal weights, the law predicted a weight of 00 for wikipedia, which we artificially set to $1\%$.
@@ D.1 Optimal domain weights
We report the optimal domain weights $h^{*}$ for the additive scaling law in Tab. 16.
Recall that  Sec. 5 defines the optimal domain weights $h^{*}(\cdot)$ as function of compute budget $(N,D)$. In the main text, we assumed uniform weights of the target domains $\mathcal{D}_{i}$. However, we can also consider a weighted average scenario, with an arbitrary weight vector $w_{i}$.
$h^{*}(w,N,D)\in\arg\min_{h\in\Delta_{k}}\sum_{i=1}^{m}w_{i}\mathcal{L}^{i}(N,D,h).$ (13)
@@
We predict the optimal domain weights for both laws, chosing each training domain $j$ as the target, that is, putting $w_{i}=1$ if $i=j$ and $w_{i}=0$ otherwise. The results are given in Fig. 7.
@@
Figure 8: The optimal domain weights are always different from the target mixture, except at the corners of the simplex. Results for a 1.3B model trained on 10B tokens, from 4 domains of The Pile.
@@
For a given $(N,D)$ pair, the optimization problem of eq. 13 defines a function $w\mapsto h^{*}(w)$ that maps the simplex onto itself. The fact that $h^{*}(w)\neq w$ indicates the surprising phenomenon that, in order to minimize the loss $\mathcal{L}^{\mathrm{ERM}}=\sum_{i=1}^{m}w_{i}\mathcal{L}^{i}(\theta)$, it is faster to instead minimize $\mathcal{L}^{*}=\sum_{i=1}^{m}h^{*}(w)_{i}\mathcal{L}^{i}(\theta)$, rather than minimizing directly $\mathcal{L}^{\mathrm{ERM}}$.
@@
To find these points, we compute the Jensen-Shannon distance, defined as $JS(w,h^{*})=\sqrt{1/2(KL(w\|m)+KL(h^{*}\|m))}$ with $m=(w+h^{*})/2$, and we look for near-zero values, in Fig. 8 and Fig. 9.
@@
In mirror of Fig. 6 we can monitor how the optimal mixture $h^{*}$ evolves as we scale the number of parameters $N$ and tokens $D$. For example, we can keep $D$ constant and scale $N$, or the opposite, or scale both simultaneously with $D\propto N$. Results are highlighted in Fig. 10. These asymptotes reach an accumulation point when $D\rightarrow\infty$ or $N\rightarrow\infty$. They depend on the speed at which $N$ and $D$ grow.
Figure 10: Asymptotes of optimal mixture when increasing $N$ and $D$ at different speeds for NMMs trained on multimodal domains. Surprisingly, there is little diffenrence between proportional scaling $\mathcal{O}(t)$ and square-root scaling $\mathcal{O}(\sqrt{t})$: both are fast enough. However, when one quantity is held constant, or only grow at logarithmic speed, the accumulation point changes.

### v2

@@ 7 Related works
Scaling laws. Scaling laws research investigates how model performance varies with training compute. Foundational studies (Hestness et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022) established that language models follow a predictable power-law relationship between performance and compute, allowing for the optimal allocation of parameters and training tokens within a specified budget. Scaling behavior has since been explored across a wide range of domains, including vision models (Fini et al., 2024; Rajasegaran et al., 2025), diffusion transformers (Liang et al., 2024), and other fields (Cheng et al., 2024; Pearce et al., 2024). While typical scaling laws consider the number of total parameters, other studies have examined the influence of both width and depth (McLeish et al., 2025), or the number of parameters allocated to the teacher and student in cse of model distillation (Busbridge et al., 2025). Sparse Mixture of Experts (MoE) models have been another focus, with investigations into how factors like sparsity, the number of experts, and routing strategies affect scaling (Krajewski et al., 2024; Clark et al., 2022; Wang et al., 2024; Abnar et al., 2025). For multimodal models, scaling laws have been explored in studies such as (Aghajanyan et al., 2023; Shukor et al., 2025). Of particular relevance is (Shukor et al., 2025), which examines native multimodal models. However, their analysis is constrained by a fixed pretraining mixture.
@@
Scaling laws for data mixtures Optimizing data mixtures for model training is a critical challenge, often requiring extensive experimentation. Recent studies have begun exploring systematic approaches to identify optimal mixtures more efficiently. For instance, Goyal et al. (2024) investigated scaling laws for data filtering in CLIP models, emphasizing data quality and repetition. Gu et al. (2024) examined scaling laws for continual pretraining of language models, predicting the optimal balance between pretraining and domain-specific data, while Bethune et al. (2025) followed the same approach, focused on forgetting in finetuning. Similarly, Chang et al. (2024) derived scaling laws that account for data quality factors such as diversity. Closer to our work, Ye et al. (2024) and Ge et al. (2024) propose scaling laws that model the loss as a function of $h$ for fixed (N, D), but they do not consider a joint law for $(N,D,h)$ as we do here. We also find that, in our experiments, for a fixed (N, D), our scaling laws extrapolate better to unseen mixtures (see appendix˜B). Further, these approaches are generally constrained to a single modality, and they consider relatively small models, below 1B parameters.
@@
Data mixture selection. The standard approach to selecting training data mixtures relies on trial and error, where different combinations are tested to determine the best-performing mixture (Soldaini et al., 2024; Lin et al., 2024; Laurençon et al., 2024; McKinzie et al., 2024; Zhang et al., 2024b; Fini et al., 2024; Shukor et al., 2023). However, this method is costly, leading to recent efforts exploring alternative strategies. Some studies adopt heuristic methods, adjusting mixture ratios based on data sizes for each domain (Rae et al., 2021; Groeneveld et al., 2024; Chung et al., 2023) or to match a target task’s distribution (Grangier et al., 2024). Others predict model performance using small models that take the mixture as input (Xie et al., 2023; Albalak et al., 2023; Liu et al., 2024; Fan et al., 2023). A third approach employs auxiliary models to rank and select high-quality training data, which has been popularized recently by large foundation models (Wenzek et al., 2019; Brown et al., 2020; Wettig et al., 2024; Penedo et al., 2024; Dubey et al., 2024).
@@ 8 Discussion
Limitations. Our current study is focused on pretraining, but continual pre-training and finetuning are also scenarios in which the mixture is important. Our scaling law predicts a generic target loss (Kaplan et al., 2020), which is known to correlate with downstream task performance (Hoffmann et al., 2022; Mayilvahanan et al., 2025). Future work may involve predicting this performance directly, like Isik et al. (2025). Furthermore, assuming no data repetition (i.e., an infinite stream of data from each domain), as is typical for LLM pretraining, is unrealistic when training with very scarce, high-quality sources. Finally, we assume that the mixture is fixed throughout training, but future works may consider a dynamic evolution of the weights (e.g., curriculum learning).
@@
Conclusion. We propose a data mixing law that predicts the loss on an arbitrary target domain, from both mixture coefficients and the compute budget $(N,D)$. Our laws hold for language, multimodal, and vision pretraining. The optimal mixture coefficients found from a small scale can be used for much larger models and training budgets, demonstrating significant improvement over domain weights found by naive grid search. This work paves the way to a principled theory of data mixture selection.
9 Acknowledgement
We thank Victor Guilherme Turrisi da Costa for technical support. We thank Marco Cuturi, Jason Ramapuram, Denise Hui, Samy Bengio, and the MLR team at Apple for their support throughout the project.
@@ Appendix A Implementation details
Optimizer Fully decoupled AdamW (Loshchilov and Hutter, 2017)
@@ A.1 LLM pretraining
For the main experiments, we use LLAMA style architectures. For the analyses with the Pile domains, we borrow architectures from (Brown et al., 2020). We use a fixed depth of 24 and change the latent dimension of the network to obtain different model scales. All hyperparameters are described in table˜5 and table˜7.
@@ A.2 Multimodal pretraining
We closely follow the implementation of (Shukor et al., 2025) and present in table˜6 the pre-training hyperparameters for the model configurations used in our scaling laws study. Training is conducted in the L3M code base (El-Nouby* et al., 2025). Our models range from 100M to 8B parameters, with width scaling accordingly while maintaining a constant depth of 24 layers. We use causal attention for text tokens and bidirectional attention for image tokens. Learning rates are adjusted based on model size, generally decreasing for larger models. These values were determined through empirical testing. Optimization is handled using a fully decoupled AdamW optimizer with momentum values set to $\beta\_1=0.9$, $\beta\_2=0.95$, and a weight decay of $1\times 10^{-4}$. Each training batch consists of 2,000 samples, totaling 2 million tokens with a 1,024-token context length. Gradients are clipped at 1.0, and training begins with a warmup phase of 1,000 iterations, followed by a constant learning rate schedule to reduce the number of experiments.
For vision inputs, we process images as $(14,14)$ patches with augmentations including Random Resized Crop (224px, scale range of [0.4, 1.0]) and Random Horizontal Flip with a 50% chance. Model training benefits from efficiency techniques such as bfloat16 precision, Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023), activation checkpointing, gradient accumulation, and sequence packing to minimize padding in the image-captioning dataset.
We assess model performance on three held-out data subsets: interleaved data (Obelics), image-caption data (HQITP), and text-only data (DCLM), following prior works (Hoffmann et al., 2022; Aghajanyan et al., 2023; Abnar et al., 2025). This setup provides a robust evaluation of model generalization across diverse data types.
@@
Optimizer Fully decoupled AdamW (Loshchilov and Hutter, 2017)
@@
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@ Appendix B Detailed extrapolation results
Table 8: Full results of experiments in figure˜3 for the LLM experiment.
@@
Table 9: Full results of experiments in figure˜3 for the multimodal experiment.
@@
Table 10: Full results of experiments in figure˜3 for the LVM experiment.
@@
Table 11: Full results of experiments in section˜6 for the LLM experiment.
@@
Table 12: Full results of experiments in figure˜3 for the cosine schedule experiment.
B.1 Comparison to the laws of Ye et al. (2024) and Liu et al. (2024)
Ye et al. (2024) propose four laws to model the behavior of the loss on a domain as a function of $h$ only, that is, for a fixed $N,D$ budget. They propose the following formulas, rewritten with notations consistent with our notation:
@@
where the parameters of the law are the $E,C,C_{i},\gamma_{i}$. Liu et al. (2024) propose to model the loss as a linear function of $h$. We compare this with the form of our scaling law equation˜2.4 when $N$ and $D$ are fixed:
@@
We fit all those scaling laws on the LLM training data, keeping only one value for $N,D$ (we take $N=200m$, $D=8B$ tokens). We only keep $25$ training mixtures to fit the laws, and report the MRE on the 8425=5984-25=59 remaining in table˜13. Note that we had trouble fitting the M3 law, which is also reported to underperform in (Ye et al., 2024). Overall, our formula gives systematically better training errors, and it most of the time translates to better testing errors on unseen mixtures. We stress once again that one of the core contributions of our work is to explain how scale interacts with data mixtures, by proposing scaling laws that take $N,D$, and $h$ as inputs. (Ye et al., 2024) only consider scaling laws with respect to $h$.
@@
Regmix PG-19 0.64 0.89
Table 13: Comparison of our scaling laws for a fixed (N, D) budget with those of (Ye et al., 2024) and (Liu et al., 2024) on the LLM data.
B.2 Comparison to Ge et al. (2024)
Ge et al. (2024) propose a scaling law to evaluate the loss the ithi-th training domains, as a function of both $h$ and number of tokens $D$:
@@
We consider a fixed size of model (N=200M) on the LLM training experiment, take only $25$ mixtures to fit the scaling laws, and report the MRE on the remaining testing set in table˜14. We see that our proposed scaling law achieves a significantly lower MRE on both train and testing data, highlighting the importance of taking all other domains into account.
@@
Table 14: Comparison of our scaling laws for a fixed model size $N$ with that of (Ge et al., 2024) on the LLM data.
Appendix C Additional analysis
C.1 Optimal domain weights
Table 15: Optimal domain weights $h^{*}$ for LLMs. Note that for the OpenHermes optimal weights, the law predicted a weight of 0 for wikipedia, which we artificially set to $1\%$.
@@ C.1 Optimal domain weights
We report the optimal domain weights $h^{*}$ for the additive scaling law in table˜15.
Recall that section˜5 defines the optimal domain weights $h^{*}(\cdot)$ as function of compute budget $(N,D)$. In the main text, we assumed uniform weights of the target domains $\mathcal{D}_{i}$. However, we can also consider a weighted average scenario, with an arbitrary weight vector $w_{i}$.
$h^{*}(w,N,D)\in\arg\min_{h\in\Delta_{k}}\sum_{i=1}^{m}w_{i}\mathcal{L}^{i}(N,D,h).$ (C.1)
@@
We predict the optimal domain weights for both laws, chosing each training domain $j$ as the target, that is, putting $w_{i}=1$ if $i=j$ and $w_{i}=0$ otherwise. The results are given in figure˜7.
@@
Figure 8: The optimal domain weights are always different from the target mixture, except at the corners of the simplex 1.3B model with 10B tokens, on 4 domains of The Pile.
@@
For a given $(N,D)$ pair, the optimization problem of equation˜C.1 defines a function $w\mapsto h^{*}(w)$ that maps the simplex onto itself. The fact that $h^{*}(w)\neq w$ indicates the surprising phenomenon that, in order to minimize the loss $\mathcal{L}^{\mathrm{ERM}}=\sum_{i=1}^{m}w_{i}\mathcal{L}^{i}(\theta)$, it is faster to instead minimize $\mathcal{L}^{*}=\sum_{i=1}^{m}h^{*}(w)_{i}\mathcal{L}^{i}(\theta)$, rather than minimizing directly $\mathcal{L}^{\mathrm{ERM}}$.
@@
To find these points, we compute the Jensen-Shannon distance, defined as $JS(w,h^{*})=\sqrt{1/2(KL(w\|m)+KL(h^{*}\|m))}$ with $m=(w+h^{*})/2$, and we look for near-zero values, in figure˜8 and figure˜9.
@@
In mirror of figure˜5 we can monitore how the optimal mixture $h^{*}$ evolves as we scale parameters $N$ and data $D$. For example, we can keep $D$ constant and scale $N$, or the opposite, or scale both simultaneously with $D\propto N$. Results are highlighted in Figure 10. These asymptotes reach an accumulation point when $D\rightarrow\infty$ or $N\rightarrow\infty$. They depend on the speed at which $N$ and $D$ grow.
Figure 10: Asymptotes of optimal mixture when increasing NN and $D$ at different speeds on multimodal data. Surprisingly, there is little diffenrence between proportional scaling $\mathcal{O}(t)$ and square-root scaling $\mathcal{O}(\sqrt{t})$: both are fast enough. However, when one quantity is held constant, or only grow at logarithmic speed, the accumulation point changes.